OpenStudy (anonymous):

Stacked cups are to be placed in a pantry. One cup is 3.25 in. high and each additional cup raises the stack 0.25 in. How many cups fit between two shelves 14 in. apart? _________ cups.

OpenStudy (anonymous):

44 cups

OpenStudy (anonymous):

3.25 + (x - 1) * .25 = 14 (x - 1) * .25 = 14 - 3.25 (x - 1) * .25 = 10.75 x - 1 = 10.75 / .25 x - 1 = 43 x = 44

OpenStudy (anonymous):

Thanks! (: I had gotten 43 when i tried. But it makes sense now :D

OpenStudy (anonymous):

Glad to help:D